OpenStudy (abbycross167):

Can someone please check the answer to my question? Please? How many solutions does the equation -2y + 3(4 – y) = -12 – 5y have? a. None b. One c. Two d. More than two

OpenStudy (kittiwitti1):

isolate y.\[-2y+3(4-y)=-12-5y\]\[-2y+12-3y=-12-5y\]\[-5y+12=-12-5y\]can you do it from here?

OpenStudy (kittiwitti1):

\[12+12=5y-5y\]\[24=0,\text{ WRONG}\rightarrow 24\ne 0\]So what do you think the answer is?

OpenStudy (abbycross167):

A?

OpenStudy (kittiwitti1):

Yes, do you know why? :P

OpenStudy (abbycross167):

because 24 =/(equal sign with the / in it like you did) 0

OpenStudy (kittiwitti1):

Yes! Good job ☺ Don't forget to close the equation :-)

OpenStudy (kittiwitti1):

I MEAN, close the question*** don't hit me > <